The amount of lemonade in a 2 litre bottle is normally distributed with a mean of 2.1 liters and a standard deviation of 0.07 liters.The portion of bottles that actually contain more than 2 liters is ?

1
Expert's answer
2022-05-03T11:32:33-0400
"P(X>2)=1-P(Z\le\dfrac{X-\mu}{\sigma})"

"=1-P(Z\le\dfrac{2-2.1}{0.07})\approx0.5568"

"55.68\%" of bottles actually contain more than 2 liters.